The measure of the supplement of an angle is 36 less than the measure of the angle. Find the measures of the angles.
Answer by cleomenius(959)

x + x - 36 = 180 degrees
2x - 36 = 180 degrees
2x = 144.
x= 72 degress
x =72 degrees. This is the smaller angle.
108 degrees, this is the larger angle.
72 + 36 = 108 degrees. This does check.
Cleomenius.